3. A bag of potato chips is packaged at . this bag of chips is transported to Denver (0.775atm) what will the new volume of the bag be at 2 degrees C ?


1
Expert's answer
2021-05-24T06:30:57-0400

Assume that a bag of potato chips is packed at standard pressure and temperature: 25 °C and 1 atm. According to the ideal gas law:

p1V1/T1 = p2V2/T2

From here:

V1/V2 = p2T1/p1T2 = (0.775 atm × 25 °C) / (1 atm × 2 °C) = 9.7

As a result, the new volume of the bag is 9.7 times smaller than the initial volume.